A survey of 25 randomly selected customers found the ages shown? (in years). The mean is 33.16 years and the standard deviation is 9.21 years.

47 35 45 21 30

42 49 29 28 35

27 29 45 35 11

34 28 33 44 39

34 25 34 32 18

a) What is the 90% confidence interval for the mean age of all customers assuming that the assumptions and conditions for the confidence interval have been met?

b) What is the margin of error?

c) How would the confidence interval change if you had assumed that the standard deviation was known to be 10.0 years?
Answer #1

a) What is the 90% confidence interval for the mean age of all customers assuming that the assumptions and conditions for the confidence interval have been met?

Mean = 33.16

SD = 9.0274

Confidence Interval is,

\bar{x}\pm z\frac{s}{\sqrt{n}}

33.16\pm 1.645*\frac{9.0274}{\sqrt{25}}

(36.1300,30.1899)

b) What is the margin of error?

Margin of Error = Critical value * Std Dev

= 1.645 * 9.0274

= 14.8500

c) How would the confidence interval change if you had assumed that the standard deviation was known to be 10.0 years?

Mean = 33.16

SD = 10

Confidence Interval is,

\bar{x}\pm z\frac{s}{\sqrt{n}}

33.16\pm 1.645*\frac{10}{\sqrt{25}}

(36.45,29.87)